OpenStudy (anonymous):

Hi can someone please help me with the following derivative: y = (sec x +3 ln tan x)^1/4.

OpenStudy (anonymous):

Ok first thing, I'm unsure of what the statement is your asking about, is it \[(\sec{x} +3 \ln({\tan{x}}) )^\frac{1}{4}\] If it is I can show you how to differentiate that if you reply. On another note I see alot of people not entering latex in an easy to read fashion (generally because they don't know how to, so I will put this here now in hope that you and others may see it. To enter latex into the box that we all use to answer or ask questions you can begin by entering it between an 'opening' and 'closing' tag. This is way better to articulate your mathematical ideas, which in mathematics needs to be done as clearly as possible because it's difficult enough already. Ok, enough rambling, the opening tag for a one or several line latex statement is this tag \[ \backslash[\] and the closing tag is this \[ \backslash ]\] So for example to write \[ (3x+3)^3 \] or this \[(3x+2)^\frac{1}{4}\] you would type EXACTLY this (minus the dollar signs) $$ \[(3x+3)^3\] or \[(3x+2)^\frac{1}{4}\] You will find now that you can be far more articulated when using latex on this site. Also It's handy having a latex symbol reference handy of which there are so many from sites to massive pdf ebooks. Ill give a basic one here http://www.artofproblemsolving.com/Wiki/index.php/LaTeX:Symbols